UT 2012-07-31
=============

CHARA/CLIMB

B1   B2   B3   B4   B5   B6
S2   E2   W1
S1   E1   W2

S1: POP 5
S2: POP 5
E1: POP 1
E2: POP 1
W1: POP 4
W2: POP 5

poor conditions, clouds moving in


*** S2-E2-W2

UT 04:15 - hd145518
   	   doing NIRO alignment
   	   recording LONG, 750 Hz, NONDEST


UT 04:40 - AS205A
   	   seeing not good (5 cm)
   	   recording LONG, 750 Hz, NONDEST
	   all fringes there, but a bit noisy (medium) due to poor conditions

UT 05:05 - hd148198
   	   recording LONG, 750 Hz, NONDEST
	   lower number of frames, since guiding on all telescopes was lost during data recording,
	   reacquired star shortly before beam sequence
	   clouds are moving in

UT 05:20 - AS205A
   	   recording LONG, 750 Hz, NONDEST, lower quality than before

UT 05:35 - hd145518
   	   recording LONG, 750 Hz, NONDEST

UT 05:50 - hd150698 (CAL for MWC863)
   	   doing NIRO alignment
   	   recording LONG, 750 Hz, NONDEST

UT 06:00 - MWC863
   	   cannot find fringes, counts too low due to clouds

UT 06:20 - hd167768
   	   recording LONG, 750 Hz, NONDEST
	   clouds, so we cannot go to target MWC297

UT 06:40 - hd177305
   	   recording LONG, 500 Hz, NONDEST

UT 06:55 - MWC614
   	   recording LONG, 500 Hz, NONDEST, good

UT 07:05 - hd177305
   	   recording LONG, 500 Hz, NONDEST

UT 07:20 - MWC614
   	   recording LONG, 500 Hz, NONDEST

UT 07:30 - hd181253
   	   recording LONG, 500 Hz, NONDEST


*** S2-E1-W2

UT 08:30 - hd194708
   	   recording LONG, 500 Hz, NONDEST

UT 08:45 - MWC340
   	   found fringes, started data acquisitions, but then clouds moved in
	   have to wait...
   	   recording LONG, 500 Hz, NONDEST
	   all fringes there, but they move a lot and are quite noisy
	   clouds moving in again... have to stop

end of night